OpenStudy (anonymous):

Let f(x) = logbase6 (4x^2-2x+6) find f'(x) So, I know that d/dx logbasea = 1/xlna so I thought it would start by looking like this 1/((4x^2-2x+6)ln6) Then from here I thought you might have to use the Quotent rule which would look like this (4x^2-2x+6)(0)-1(8x-2)/((4x^x-2x+6)ln6)^2 which the top simplified would be -8x+2/((4x^2-2x+6)ln6)^2 but, now I am stuck please see if what I am doing is correct and help me figure out what is next. thanks.

OpenStudy (zugzwang):

To be clear, \[\frac{d}{dx}\log_{6}(4x^{2}-2x+6)\] Is what you're trying to find?

OpenStudy (anonymous):

Yes

OpenStudy (zugzwang):

You did well up to \[\frac{1}{(4x^{2}-2x+6)\ln6}\] But you will multiply this to the derivative of \[4x^{2}-2x+6\] and not to the derivative of \[\frac{1}{(4x^{2}-2x+6)\ln6}\] So for this particular problem you don't have to worry about the quotient rule... :)

OpenStudy (anonymous):

so you would just do the 1/(4x^2-2x+6)ln6 *8x-2?

OpenStudy (zugzwang):

Yes, that looks about right, and you're done!

OpenStudy (anonymous):

Thanks!

OpenStudy (zugzwang):

No problem